5. Let Xi i = 1,2, . ,N be i.i.d. U(0,1). Let Z = max{X1, .,Xn} and find Fz.
Answer #1

solutions - f2 = P(253) = P(max f20,8 -- aný sa) 2 - P(X, 52,2253,-- in S2) = P(X, 53).P(9252) -- Plansa ) as 7; as lid - - -